# PQA: Zero-shot Protein Question Answering for Free-form Scientific Enquiry with Large Language Models

### v1

@@ Abstract
We introduce the novel task of zero-shot Protein Question Answering (PQA) for free-form scientific enquiry. Given a previously unseen protein sequence and a natural language question, the task is to deliver a scientifically accurate answer. This task not only supports future biological research, but could also provide a test bed for assessing the scientific precision of large language models (LLMs). We contribute the first specialized dataset for PQA model training, containing 257K protein sequences annotated with 1.97M scientific question-answer pairs. Additionally, we propose and study several novel biologically relevant benchmarks for scientific PQA. Employing two robust multi-modal architectures, we establish an initial state-of-the-art performance for PQA and reveal key performance factors through ablation studies. Our comprehensive PQA framework, named Pika, including dataset, code, model checkpoints, and a user-friendly demo, is openly accessible on github.com/EMCarrami/Pika, promoting wider research and application in the field.
Machine Learning, Multimodel Large Language Models, Protein Sequence
11footnotetext: Relation Therapeutics, London, UK22footnotetext: Ludwig Maximilian University of Munich, Germany*Corresponding author: Eli M Carrami, eli@relationrx.com, eli.carrami@gmail.com
@@ 1 Introduction
Proteins, essential to biological functions, are complex macromolecules that perform a myriad of cellular roles determined by their complex structures and interactions. As a polymer of amino acids, drawn from a pool of 20 natural ones, the sequence in which these building blocks are arranged dictates the protein’s three-dimensional structure, which is critical for its function. These structures facilitate interactions with various molecules, underpinning numerous biological processes. Given the critical role of proteins in both fundamental biology and applied biomedical research, a deep understanding of their structures and functions is crucial. Despite significant advances in deciphering proteins’ 3D structures, there remains a pressing need for innovative methodologies to facilitate the computational study of their biochemical and functional properties.
Currently, this challenge is primarily addressed using task-specific models, necessitating extensive data collection and model training for each query. For instance, submissions to the Critical Assessment of Protein Function Annotation algorithms (CAFA) aim to predict the functional annotations (e.g., GO terms) for new protein sequences, as required by the multi-year challenge (Function-SIG, 2024). Similarly, several other task-specific models have been developed to predict specific biochemical properties of proteins, such as ligand binding (Wei et al., 2022) or thermal stability (Blaabjerg et al., 2023).
In this work, we introduce the task of zero-shot Protein Question Answering (PQA) as a novel endeavor to unify protein sequence-related enquiries. To this end, we have developed the Pika framework, comprised of a curated and unbiased dataset specifically designed for the PQA domain, equipped with scientific Q/A pairs for instructional training as well as a robust benchmarking strategy to enable effective evaluation of PQA models. Alongside, we also propose multimodal large language models as a potential solution to this task. Circumventing the need for custom-designed models or datasets for each distinct task, this approach allows for unconstrained free-form scientific enquiry and open-ended answer generation by the large language model (LLM). This is particularly timely, given the significant strides in natural language processing that have led to the development of powerful LLMs. These models, besides their capability in processing natural questions, also encapsulate a large body of knowledge which could provide further context for the model to provide accurate and coherent responses. This paradigm shift enables a more flexible and efficient exploration of protein functionalities, bypassing the need for extensive model training and data collection for individual queries.
In particular, we train and evaluate two multimodal protein-text architectures combining the ESM2 (Lin et al., 2022) protein language model (PLM) with the Phi-2 LLM (Microsoft, 2023), showcasing the seamless integration of protein sequence analysis with natural language processing. This approach yields accurate responses to scientific enquiries regarding the biochemical and functional properties of previously unseen proteins based purely on their sequences, representing an important step forward in this domain. It is also essential to highlight that, our research, leveraging the robust yet modestly scaled Phi-2 LLM, serves as a robust proof-of-concept to demonstrate the potential of this approach. Our results suggest that the strategic adoption of this methodology, especially with more advanced LLMs, has the potential to challenge the current state-of-the-art in task-specific models. Consequently, our approach, by facilitating scientific enquiries through natural language question answering, sets the stage for broader accessibility and efficiency in future research endeavors. This development promises to expedite the discovery process and hypothesis generation, thus reshaping the landscape of biological and biomedical research.
2 Related Work
2.1 Multimodal Language Models
Recent advancements in Multimodal Language Models have significantly enriched the interaction between text and other modalities. Central to this progress are models like SimVLM (Wang et al., 2021), Frozen (Tsimpoukelli et al., 2021), Flamingo (Alayrac et al., 2022), PaLI (Chen et al., 2022) for visual content, alongside innovations like AudioLM (Borsos et al., 2023) for audio data. These models blend the strengths of vision encoders built by convolutional neural networks (LeCun et al., 2015) or transformers (Vaswani et al., 2017; Dosovitskiy et al., 2020), with language models to overcome the limitations of traditional single-modality models. Furthermore, the scope of multimodal models can extend beyond vision, audio, and video to include agent actions (Reed et al., 2022), biomedical data (Zhang et al., 2023; Tu et al., 2023) or molecules (Edwards et al., 2022).
2.2 Protein Language Models
This work draws upon a large body of previous research in creating rich representations of proteins for 3D structure prediction or functional classification (ESM family (Rives et al., 2021; Lin et al., 2022) and AlphaFold (Jumper et al., 2021)). The ESM2 family of PLMs have been trained on a large dataset of protein sequences in a self-supervised manner to yield rich representations for protein sequences through learning the “language of proteins”. These models have demonstrated strong performance in various downstream tasks, including protein folding, functional prediction, and disease-variant effect prediction, highlighting their rich and representative embedding space.
The previous work on integrating text with protein sequences has primarily focused on improving the protein representation by leveraging a large body of unstructured biomedical text or knowledge graphs (ProtST (Xu et al., 2023), OntoProtein (Zhang et al., 2022)), with a focus on functional classification or protein sequence generation from textual descriptions. Conversely, ProteinChat (Guo et al., 2023) has utilized LLM as a knowledge retrieval platform to facilitate discussions about 3D structure databases. These efforts have also led to the creation of two datasets, ProtDescribe (Xu et al., 2023) and PDB-QA (Wang et al., 2023). While PQA has been explored in previous and concurrent research, these have primarily been limited to evaluating knowledge retrieval due to the limitations of available datasets and evaluation metrics (details in Section 3.1). As a result, the task of conducting free-form zero-shot scientific enquiry of new protein sequences remains unexplored.
3 PQA: Dataset & Benchmark
In this section, we detail our dataset and benchmark designs for the Pika framework. Starting with identifying limitations of existing protein-text databases, we then outline the essential criteria for a dedicated PQA dataset, followed by a description of the creation process and the quality control of our specialized dataset, Pika-DS. Finally, we introduce several biochemically motivated benchmarks, essential for assessing the scientific accuracy of PQA models.
3.1 Specialized PQA Dataset
@@
The scientific PQA task is aimed at delivering accurate responses to free-form questions based on an unseen protein sequence. This task emerges from the need for scientific exploration of protein functions via natural language question answering. Training and evaluation of multimodal models for the PQA task necessitates comprehensive datasets with scientific textual annotations linked to corresponding protein sequences. Currently two PQA-relevant datasets are available. While these datasets are suitable for their intended tasks, we found them unfit for the specific purpose of scientific PQA for reasons described below:
@@
    ProtDescribe dataset: This dataset (Xu et al., 2023) contains textual annotations from three fields of uncurated SwissProt entries, risking bias and data leakage due to high sequence similarity and over-representation of certain protein families. Its use of fields that are not directly inferable from protein sequence (e.g., cellular interaction) limits its utility for PQA. Conversely, it misses out on features like catalytic reactions which are often subject of scientific queries. Crucially, lacking Q/As, it fails to directly support PQA assessments. Therefore, while valuable for enhancing protein representations and text-to-protein generation, as intended by the authors, this dataset exhibits significant limitations for the scientific PQA task.
@@
    PDB-QA dataset: (Guo et al., 2023) With 30 predefined Q/As about 3D structures from the PDB database, while suitable for knowledge retrieval tasks, PDB-QA’s focus on entry-specific details (e.g., submission date, analysis software) renders it ineffective in the context of scientific PQA.
3.2 Dataset Creation
To address the gaps discussed in the previous subsection, we created Pika-DS, the first specialized and unbiased PQA dataset, accompanied with respective biologically relevant benchmarks for model training and evaluation. We designed this dataset to fulfill three key criteria essential for the training and benchmarking of PQA models:
@@ 3.2 Dataset Creation
  • It includes pertinent benchmarking metrics to assess model performance especially in zero-shot settings.
# total # unique # words # tokens
Proteins 257 K 252 K - 105.3 M
Info fields 1.17 M 319 K 10.8 M 29.5 M
Summaries 1.33 M 861 K 22.5 M 36.4 M
Questions 1.71 M 98 K 15.7 M 14.9 M
Answers 1.71 M 783 K 18.4 M 31.8 M
Table 1: Pika-DS size and token counts. Total and Unique values for Summaries represent the number of sentences.
Following these criteria, Pika-DS was prepared in three phases:
@@
    UniProt also offers similarity-based clustering of its entries at various sequence similarity thresholds of 100%, 90%, and 50%, known as UniRef clusters. For instance, protein sequences that share a minimum of 50% pairwise sequence similarity will be assigned the same UniRef50 cluster. Recognizing a strong bias towards more commonly studied protein families in SwissProt (Figure B.1), where nearly 50% of the SwissProt sequences belong to the largest 6% of clusters, we limited Pika-DS to a maximum of two most informative sequences per UniRef50 cluster using a custom algorithm (Section A.2). This resulted in a debiased set of over 257,000 protein sequences and 1.17 million information fields.
  3. 3.
    Lastly, we instructed GPT3.5 to process the information fields for each protein entry to create the Pika-DS’s three main components (example: Table B.1):
    • QAs: Several diverse Q/A pairs for each information field, formatted for LLM training.
    • Metrics: Single-word answers to a set of predefined scientific questions serving as the ground-truth for our Biochem-Lite benchmarks (Section 3.4).
    The prompts for all instructions were fine-tuned to ensure GPT3.5’s adherence to using only the given data, summarizing complex biological reactions accurately, and avoiding speculative or ambiguous language. Examples in Section A.3.
The final Pika-DS comprises 257,167 protein sequences, carefully selected from 185,128 UniRef50 clusters. This dataset is enriched with detailed descriptors for each sequence: a summary statement divided into sentences, multiple scientific Q/A pairs, and the ground-truth answers for our predefined benchmarking questions. This dataset encapsulates over 105 million protein sequence tokens, approximately 36.4 million tokens in textual summaries, and nearly 47 million tokens in Q/A pairs (Table 1).
3.3 Dataset Analysis
In this subsection, we analyze various facets of our finalized PQA dataset to validate its quality and suitability for training and evaluating multimodal models for the purpose of open-ended scientific enquiries of protein sequences.
Protein over-representation Analysis:
We used pre-computed protein embeddings from UniProt database and visualized them using UMAP (Uniform Manifold Approximation and Projection for Dimension Reduction(McInnes et al., 2018)), highlighting the entries that belong to the top 100 largest clusters. Visual comparison of distribution of protein embeddings before and after our Uniref50-based filtering confirms a strong reduction of over-represented protein families in Pika-DS while maintaining the rich diversity of the dataset (Figure B.2).
Token Count Analysis:
In creating the Q/A pairs using GPT3.5, we instructed the model to cover a broad spectrum of queries and to produce detailed answers suitable for training other LLMs, anticipating elaborate responses rather than single-word answers. This was confirmed by our analysis of token counts in the summary sentences, questions, and answers of the final dataset. This revealed that while questions are typically shorter (8.5 tokens on average), the summary sentences and answers are longer and more extensive (on average 27 and 16.5 tokens). This presents an adequate number of tokens per example, providing sufficient context for the training of multimodal PQA models (Figure 2a).
Word Frequency Analysis:
To ensure the diversity of content in summary statements as well as Q/A pairs, we examined the frequency of words in the initial five positions of each of these categories. As anticipated, common words such as “the” and “it”, in summaries, and interrogatives like “What” and “How” in questions, dominated the first three positions. Similarly, answers often began with “the” or “this” as well as “yes/no” structures, followed by the word “protein”. However, the remaining positions demonstrated a significant lexical diversity. This observation confirmed that our dataset does not exhibit a substantial bias, offering the necessary diversity for effective model training (Figure 2b).
These analyses underscores the high quality and diversity of the Pika-DS, affirming its suitability for further research in this domain.
Metric ID Question Example MLP task
label Self-Pika’s response
mw MALE

### v2

# PQA: Zero-shot Protein Question Answering for Free-form Scientific Enquiry with Language Models

@@ Abstract
Understanding protein structure and function is crucial in biology. However, current computational methods are often task-specific and resource-intensive. To address this, we propose zero-shot Protein Question Answering (PQA), a task designed to answer a wide range of protein-related queries without task-specific training. The success of PQA hinges on high-quality datasets and robust evaluation strategies, both of which are lacking in current research. Existing datasets suffer from biases, noise, and lack of evolutionary context, while current evaluation methods fail to accurately assess model performance. We introduce the Pika framework to overcome these limitations. Pika comprises a curated, debiased dataset tailored for PQA and a biochemically relevant benchmarking strategy. We also propose multimodal large language models as a strong baseline for PQA, leveraging their natural language processing and knowledge. This approach promises a more flexible and efficient way to explore protein properties, advancing protein research. Our comprehensive PQA framework, Pika, including dataset, code, and model checkpoints, is openly accessible on Github, promoting wider research in the field.
@@ 1 Introduction
Proteins, essential to biological functions, are complex macromolecules that perform a myriad of cellular roles determined by their complex structures and interactions. As a polymer of amino acids, drawn from a pool of 20 natural ones, the sequence in which these building blocks are arranged dictates the protein’s three-dimensional structure, which is critical for its function. Given the critical role of proteins in both fundamental biology and applied biomedical research, a deep understanding of their structures and functions is crucial. Despite significant advances in deciphering proteins’ 3D structures, there remains a pressing need for innovative methodologies to facilitate the computational study of their biochemical and functional properties.
Currently, training individual models tailored to specific tasks are the primary approach to computationally studying the biochemical and functional properties of proteins, requiring extensive data collection and training for each unique task. For instance, submissions to the Critical Assessment of Protein Function Annotation algorithms (CAFA) aim to predict the functional annotations such as GO terms for new protein sequences, as required by the multi-year challenge (Function-SIG, 2024). Similarly, several other task-specific models have been developed to predict specific biochemical properties of proteins, such as ligand binding (Wei et al., 2022) or thermal stability (Blaabjerg et al., 2023). To address the limitations of current approaches, we propose to unify protein sequence-related enquiries under a more generic task of zero-shot Protein Question Answering (PQA), where there can be free-form inquiries about known or novel protein sequences.
The development and assessment of effective zero-shot PQA models critically depend on the availability of high-quality datasets and robust evaluation strategies, both of which present significant challenges in the current research landscape; unfortunately, existing datasets available for related tasks often exhibit significant limitations. They frequently demonstrate biases towards specific protein families, overlook the critical evolutionary relationships between proteins, and may contain noisy or unreliable annotations. These shortcomings hinder the development of effective PQA models that can generalize across diverse proteins and accurately answer complex queries.
Similarly, the robust evaluation of PQA models necessitates biochemically-relevant benchmarking strategies that focus on the scientific accuracy of model predictions within the context of the specific questions posed. Previous research efforts have not adequately addressed these challenges. They often neglect the impact of evolutionary relationships on information leakage, which can lead to overestimation of model performance. Moreover, they predominantly rely on metrics like BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004), which have been consistently shown to be inadequate for assessing the accuracy of scientific statements (Mathur et al., 2020). This lack of suitable evaluation tools further impedes progress in the field.
To address these limitations, we have developed the Pika framework, comprised of a curated and debiased dataset specifically designed for the PQA domain, equipped with scientific question and answer (QA) pairs for instructional training as well as a robust and biochemically relevant benchmarking strategy to enable effective evaluation of PQA models. Alongside, we also propose multimodal large language models (LLMs) as a potential solution to PQA and create evolution-aware splits for assessment of their performance. Besides their capability in processing natural questions, LLMs encapsulate a large body of knowledge which could provide further context for the model, enabling it to perform a more flexible and efficient exploration of protein functionalities, bypassing the need for extensive model training and data collection for individual queries.
Finally, comparing the performance of multimodal LLMs with various relevant baselines on our evolution-aware data splits, we show that these models are a promising direction for PQA. In particular, we train and evaluate two multimodal protein-text architectures combining the ESM2 (Lin et al., 2022) protein language model (PLM) with the Phi-2 LLM (Microsoft, 2023), showcasing the seamless integration of protein sequence analysis with natural language processing while highlight the shortcomings of these models in particular when dealing with evolutionary distant proteins. Our results suggest that the strategic adoption of this methodology, especially with more advanced LLMs, has the potential to challenge the current state-of-the-art in task-specific models. It is essential to highlight that, our research, leveraging the robust yet modestly scaled Phi-2 LLM, serves as a robust proof-of-concept to demonstrate the potential of this approach. We hope our work fuels further research and opens new avenues towards performant PQA models.
2 Related Work
Since PQA is distinct from other question-answering tasks, existing scientific QA datasets such as ScienceQA (Lu et al., 2022), PubMedQA (Jin et al., 2019) or SQuAD (Rajpurkar et al., 2016, 2018) are not suitable for PQA training or evaluation. This is because unlike standard QA tasks where the answer to the question can be inferred using logic or existing knowledge, in PQA the answer to the question must be extracted from cross-modality embeddings via the LLM, therefore necessitating specialized protein-text datasets that annotate protein sequences with relevant labels.
Previous studies on integrating text and protein sequences have primarily leveraged large-scale unstructured biomedical text or knowledge graphs to refine protein representations, facilitating downstream tasks like functional classification and sequence generation (e.g. ProtST (Xu et al., 2023), OntoProtein (Zhang et al., 2022)). Alongside, ProteinChat (Guo et al., 2023) has utilized LLMs for knowledge retrieval to facilitate discussions about 3D structure databases, and concurrent to our work, Mol-Instructions (Fang et al., 2024) focuses on improving LLM’s understanding of biomolecules (e.g., proteins) by fine-tuning existing LLMs. These recent efforts while introducing valuable protein-text datasets have limitations that make them unsuitable for scientific PQA, as detailed below:
  1. 1.
    ProtDescribe (Xu et al., 2023): This dataset contains textual annotations from three fields of uncurated SwissProt entries, risking bias and data leakage due to high sequence similarity and over-representation of certain protein families (Fig. B.1). Also, its use of fields that are not directly inferable from protein sequence alone (e.g., cellular interaction) limits its utility for PQA. Conversely, it misses out on features like catalytic reactions which are often subject of scientific queries. Crucially, lacking QAs, it fails to directly support PQA assessments. Therefore, while valuable for enhancing protein representations and text-to-protein generation, as intended by the authors, this dataset is not suitable for the scientific PQA.
  2. 2.
    PDB-QA (Guo et al., 2023) : With a limited set of 30 predefined QAs on 3D structures from the PDB database, while suitable for knowledge retrieval tasks, its focus on entry-specific details (e.g., submission date, analysis software) renders it ineffective in the context of scientific PQA.
  3. 3.
    Mol-Instructions (Fang et al., 2024): This dataset was released concurrent to our work. The protein section of Mol-Instructions represents a relevant set of template-based textual annotations for proteins tailored towards five downstream tasks, including question-answering. The dataset is curated from SwissProt and is debiased using a 90% similarity threshold. This lenient threshold results in an abundance of highly related proteins, which leads to bias and in the absence of an evolution-aware splitting strategy could cause leakage across data splits.
Furthermore, in all previous and concurrent research authors have relied on BLEU or ROUGE metrics for evaluating the scientific accuracy of generate captions or responses. However, these metrics are not suitable for assessing accuracy (Mathur et al., 2020), therefore limiting the scientific scope of past research in PQA domain. As a result, due to the lack of relevant benchmarking strategies as well as biases and noise in existing datasets, the task of free-form zero-shot scientific enquiry of new protein sequences remains unexplored.
@@ 2 Related Work
3 Pika Framework
Here we detail our dataset, benchmarking and baseline designs for the Pika framework (Fig. 1). All baselines, model architectures and benchmarking were implemented in PyTorch Lightning , and the complete codebase is accessible on github.com/EMCarrami/Pika.
3.1 Pika Dataset
@@
The scientific PQA task is aimed at delivering accurate responses to free-form questions based on an unseen protein sequence. This task emerges from the need for scientific exploration of protein functions via natural language question answering. Therefore, the training and evaluation of multimodal models for the PQA task necessitates comprehensive datasets with scientific textual annotations linked to corresponding protein sequences. We deemed the following three criteria as essential for a specialized PQA dataset:
@@
  3. 3.
    Supports relevant benchmarking to assess model performance especially in zero-shot settings.
Since, as discussed in section 2, none of the existing datasets meet all these criteria we created Pika-DS (See sections A.1, A.2, A.3 for details), the first specialized and debiased PQA dataset, accompanied with respective biologically relevant benchmarks for model training and evaluation (summary statistics in Table B.1 and example in Table B.2).
Briefly, we gathered all SwissProt entries from UniProt database (Consortium, 2022) and extracted an expert-curated list of scientific information fields covering a wide range of relevant properties. We then debiased the sequences using a strict 50% similarity threshold. Finally, we employed GPT3.5 API to process the information fields for each protein entry using systematically optimized prompts (Sections A.4 and A.5) to create the Pika-DS’s three main components:
  • QAs: Several diverse QA pairs for each information field, formatted for LLM training.
  • Metrics: Single-word answers to a set of predefined scientific questions serving as the ground-truth for our Biochem-Lite benchmarks (Section 3.2.2).
3.1.1 Pika-DS Quality Control
The final Pika-DS comprises 257,167 protein sequences, selected from 185,128 UniRef50 clusters. This dataset is enriched with detailed descriptors for each sequence: a summary statement divided into sentences, multiple scientific QA pairs, and the ground-truth answers for our predefined benchmarking questions. This dataset encapsulates over 105 million protein sequence tokens, approximately 36.4 million tokens in textual summaries, and nearly 47 million tokens in QA pairs (Table B.1). To ensure the high quality of Pika-DS and suitability of generated textual annotations for training multimodal LLMs we assessed various aspects of our dataset.
Human evaluation of GPT3.5 generated annotations:
Given the potential limitations of GPT3.5 generated annotations (e.g., hallucination), we conducted a thorough human evaluation of the Pika-DS. For this, 100 randomly selected examples from Pika-DS were manually evaluated by an expert biochemist. Our analysis revealed that out of a total of 1204 summary sentences and QA pairs, 5 were incorrect (0.4%), 21 were of poor quality (1.7%), and 43 were irrelevant (3.6%), resulting in over 94% of the annotations being correct, relevant, and of high quality. Furthermore, only 5 out of 660 metric ground truths were found to be incorrect, yielding a 99.2% accuracy in ground truth metric values. Furthermore, consultations with two expert biologists confirmed that the QA pairs generated by GPT-3.5 in Pika-DS are relevant and scientifically sound, closely matching what they would derive from the provided input fields. They also unanimously agreed that GPT-3.5 questions offer greater diversity compared to their potential queries. These results underscore the effectiveness of our prompt optimization approach for generating high-quality labels using GPT3.5 based on SwissProt information fields and is consistent with a recent study that has found LLM-generated captions can exhibit higher diversity than those created by humans, leading to enhanced training of multimodal LLMs (Sharifzadeh et al., 2024).
Token Count Analysis:
In creating the QA pairs using GPT3.5, we instructed the model to cover a broad spectrum of queries and to produce detailed answers suitable for training other LLMs, anticipating elaborate rather than single-word answers. This was confirmed by our analysis of token counts in the summary sentences, questions, and answers of the final dataset. This revealed that while questions are typically shorter (8.5 tokens on average), the summary sentences and answers are longer and more extensive (on average 27 and 16.5 tokens). This presents an adequate number of tokens per example, providing sufficient context for the training of multimodal PQA models (Figure 2a).
Word Frequency Analysis:
To ensure the diversity of content in summary statements as well as QA pairs, we examined the frequency of words in the initial five positions of each of these categories. As anticipated, common words such as "the" and "it", in summaries, and interrogatives like "What" and "How" in questions, dominated the first three positions. Similarly, answers often began with "the" or "this" as well as "yes/no" structures, followed by "protein". However, the remaining positions demonstrated a significant lexical diversity. This observation confirmed that our dataset does not exhibit a substantial bias, offering the necessary diversity for effective model training (Figure 2b).
Protein over-representation Analysis:
We used pre-computed protein embeddings from UniProt database and visualized them using UMAP (McInnes et al., 2018), highlighting the entries that belong to the top 100 largest clusters. Visual comparison of distribution of protein embeddings before and after our Uniref50-based filtering confirms a strong reduction of over-represented protein families in Pika-DS while maintaining the rich diversity of the dataset (Figure B.2).
3.2 Pika Benchmarks
3.2.1 Motivation & Design Criteria
@@
In consultation with domain experts, we identified five core protein properties at distinct difficulty levels that form the basis for our scientific benchmarking questions: molecular weight (mw), co-factor binding, sub-cellular localization, protein domains, and enzymatic reaction.
Alongside these benchmarking questions, we also require a robust strategy to assess the correctness of open-ended answers to them. Four possible assessment strategies could be imagined: (1) Exact matching of statements, (2) Keyword comparison, (3) Comparison using stronger LLMs (xLLMs), and (4) Human evaluation. While exact statement matching is useful in some research domains, it lacks the rigour required for assessing scientific accuracy. On the other hand, human evaluation for specialized domains is extremely costly at large scales. However, we can use human supervision to ensure the quality of comparisons conducted by xLLMs. As a result we selected keyword comparison and use of xLLMs guided by human experts for Pika framework.
3.2.2 Two-Tiered Benchmarking System
The use of xLLMs for response evaluation is more accurate but computationally prohibitive as compared to keyword comparison, which pragmatically balances scientific accuracy with computational efficiency. This balance of performance vs efficiency motivated us to design a two-tiered benchmarking system comprised of a light-weight benchmark employing keyword comparison for efficiency, and a rigorous xLLM-based benchmark for scientific fidelity.
Metric ID Question Example Label Example Response
mw MALE $\downarrow$